OpenStudy (anonymous):

HEEEEELLLLLLLPPPPPPPP! ILL MEDAL! The angle made by the ladder with the ground is_________ degrees, and the length of the ladder is_________ inches

OpenStudy (mathmale):

Your illustration, please?

OpenStudy (anonymous):

A ladder is placed 30 inches from a wall. It touches the wall at a height of 50 inches from the ground sorry I forgot to add this...

OpenStudy (anonymous):

@nora1205 - could you please draw a diagram for the question as understood by you ? it would help in understanding your approach...

OpenStudy (jsizzie101):

what do you need to find?

OpenStudy (anonymous):

to be honest I have no idea, that's all it gave me A ladder is placed 30 inches from a wall. It touches the wall at a height of 50 inches from the ground The angle made by the ladder with the ground is_________ degrees, and the length of the ladder is_________ inches

OpenStudy (anonymous):

it didn't give me a picture or anything

OpenStudy (mathmale):

Draw a vertical wall, draw the ground, and draw a ladder. I'm sure you could do that.

OpenStudy (anonymous):

 yeah I actually can do that :)

OpenStudy (anonymous):

\[\tan ^{-1}\left(\frac{50}{30}\right)=59.0362 {}^{\circ} \]\[\sqrt{50^2+30^2}=58.3095 \]

OpenStudy (anonymous):

thank you for the help,